OpenStudy (anonymous):

Simplify 4 x 5 – 4 2 x 6

OpenStudy (anonymous):

4/2

OpenStudy (unklerhaukus):

\[4 \times 5 – 4 \times( 2 \times 6)\]?

OpenStudy (ajprincess):

can u plzz retype ur question again?

OpenStudy (anonymous):

4 x 5 - 4 / 2 x 6

OpenStudy (ajprincess):

Do u knw BODMAS rule?

OpenStudy (unklerhaukus):

\[4 \times 5 - \frac{4 }{ 2\times 6}\]\[=20-\frac{4}{12}\]\[=20-\frac13\]\[=\frac{60}{3}-\frac13\]\[=\frac{59}{3}\]

OpenStudy (unklerhaukus):

i still think im reading the original equation wrong

OpenStudy (unklerhaukus):

what is the numerator (the bit on top of the of the fraction)

OpenStudy (anonymous):

dont worry i got hte answer all ready

OpenStudy (precal):

@ajprincess what is BODMAS rules? just curious

OpenStudy (ajprincess):

B- bracket O-of D-division M-multiplication A-addition S-Subtraction When there are multiple operations u need to perform them in this order.

OpenStudy (precal):

what is o -of we use PEDMAS Parentheis (inside) Exponents Multiplication/Division Left to right Addition/Subtration left to right